OpenStudy (anonymous):

Solve the equation sin(2t) + cost= ¼ on the interval [0, 2π]. Express answer(s) to three decimal places.

OpenStudy (anonymous):

I graphed y=sin(2t)+cost and then y=1/4, but they are coming out as straight lines. did I do something wrong?

OpenStudy (anonymous):

\[ \sin(2t)+\cos(t)={1\over4}\\ 2\sin(t)\cos(t)+\cos(t)={1\over4}\\ \cos(t)\left[2\sin(t)+1\right]={1\over4}\\ \cos(t)={1\over 4\left[2\sin(t)+1\right]}\\ \;\\ \text{we have the relation}\\ \sin^2(t)+\cos^2(t)=1\\ \sin^2(t)+{1\over16\left[2\sin(t)+1\right]^2}=1 \] solve the last quadratic equation for t